OpenStudy (lolaismy):

An 80.0-gram sample of a gas was heated from 25 °C to 225 °C. During this process, 346 J of work was done by the system and its internal energy increased by 8625 J. What is the specific heat of the gas? I keep getting .517J but the homework online is telling me I'm wrong. I really need to get this answer and understand where I went wrong! Thank you in advance!

OpenStudy (australopithecus):

By the first law of thermodynamics: \[\Delta U = q + w\] q is heat energy w is work energy U is the change in internal energy Note: Work done by the system is negative Work done on the system is positive The question gives you two of the variables so you can solve for q Now we need the equation for heat capacity \[C = \frac{q}{m*\Delta T}\] where, m = mass in grams q = heat energy C = cpecific heat capacity \[\Delta T = T_{final\ Temperature} - T_{initial\ Temperature} = Change\ In\ Temperature\] Solve for C and you are done

OpenStudy (australopithecus):

Remember temperature is in kelvin but it doesnt really matter since the 273 will cancel out anyways when you find the change

OpenStudy (lolaismy):

@Australopithecus So I solved the way you demonstrated using my values and I got a new calculation of .4629 J... Does this value seem correct?

OpenStudy (australopithecus):

Did you read everything I wrote?

OpenStudy (australopithecus):

I think you made a mistake

OpenStudy (australopithecus):

Show me your work? I can tell you where you are making a mistake

OpenStudy (lolaismy):

@Australopithecus So I tried again and I got the same answer as I did in the beginning which according to my online homework is wrong. Using equation #1.) 200degrees=q+8971J q+-8771J Equation #2.) C=-8771/((80)(9200degrees)) = -.548187 J What am I doing wrong?! ):

OpenStudy (joannablackwelder):

deltaU is the change in internal energy, so \[\Delta U = q + w\] \[8625=q+346\]

OpenStudy (joannablackwelder):

Does that step make sense?

OpenStudy (joannablackwelder):

Because it says that the change in internal energy is 8625 J and the work done is 346 J

OpenStudy (lolaismy):

@JoannaBlackwelder okay so following through with that, I get that q=8279 then if i place that into the other equation 8279/((80grams)(200deltaT)) then I get that c=.517 which is as well wrong in accordance to my online homework!

OpenStudy (joannablackwelder):

You put before that the unit is J, which is incorrect. What unit did you use?

OpenStudy (australopithecus):

It is work done by the system so work is negative

OpenStudy (lolaismy):

(J/g*C)=.517

OpenStudy (joannablackwelder):

Oh, and I just realized that it says that work was done by the system.

OpenStudy (australopithecus):

that is his mistake I believe

OpenStudy (joannablackwelder):

Yep, I think you are right.

OpenStudy (australopithecus):

8625 = q + (-346) q = 8625 + 346

OpenStudy (australopithecus):

The mistake was in calculating q

OpenStudy (lolaismy):

so then q=8971, 8971/((80)(200))=C, C=.5606 (J/g*C)

OpenStudy (lolaismy):

or should that be a negative value because it is work done by the system?

OpenStudy (australopithecus):

it would be Joules/K*g

OpenStudy (australopithecus):

Remember when doing chemistry questions always convert to Kelvin unless otherwise told

OpenStudy (australopithecus):

but in this case it wont matter because 273 is canceled out but it is still converted T = (225+273)K - (25 + 273)K

OpenStudy (australopithecus):

= 200K

OpenStudy (lolaismy):

The question wants it in (J/g*C) But im not sure if that makes a difference in this question

OpenStudy (australopithecus):

oh then it will be the same

OpenStudy (lolaismy):

so then the final answer would be (J/g*C)=-.56068 (negative not positive)?

OpenStudy (australopithecus):

you shouldn't be getting a negative

OpenStudy (lolaismy):

oh yeah you're right I'm just getting positive .56068

OpenStudy (australopithecus):

that is what I got

OpenStudy (lolaismy):

AND THE ANSWER IS CORRECT!!! Thank you so much for breaking it down for me!!! :D

OpenStudy (australopithecus):

No problem sorry for taking so long to get back to you

OpenStudy (lolaismy):

It's fine I appreciate your time ^^